OpenStudy (blurryface):

Solve the equation using the Zero-Product Property. -8n(8n-8)=0

OpenStudy (blurryface):

@Sachintha @nincompoop

OpenStudy (blurryface):

@Kainui @UsukiDoll

OpenStudy (kainui):

if a*b=0 then either a=0 or b=0

OpenStudy (blurryface):

I know how to use the Zero-Product Property, but I'm unsure what to do to get the equation in form if ab=0 then a=0 or b=0

OpenStudy (kainui):

make a guess

OpenStudy (blurryface):

umm use the distribute property?

OpenStudy (blurryface):

(-8n*8n)(-8n*-8)=0 Is this correct?

OpenStudy (blurryface):

hello?

OpenStudy (kainui):

Well, if you understand how to use the zero product property, can you explain it? That might be helpful.

OpenStudy (kainui):

Right now you have -8n(8n-8)=0. Do you agree that -8n is being multiplied by 8n-8?

OpenStudy (blurryface):

yea

OpenStudy (blurryface):

this is bs, thanks for your "help".. smh

OpenStudy (kmeis002):

Kainui did a fine job leading you to the answer. Look at what you know: 1) You agree that if \( ab = 0 \implies a = 0 \) or \( b = 0 \). 2) You agree that \(-8n(8n-8) = 0 \) is \( -8n \) being multiplied by \( (8n-8) \) Since \(-8n(8n-8) \) equals \( 0 \) AND it is the multiplication of two factors, it would follow that you can use the Zero-Product Property. The question is what is \(a\) and what is \(b\) in \(-8n(8n-8) = 0 \)?

OpenStudy (blurryface):

that's what I need help on.. duhh

OpenStudy (kmeis002):

You agree that \(-8n\) and \((8n-8)\) are multiplying by each other. What are \(a\) and \(b\) doing?

OpenStudy (blurryface):

(-8n*8n)= a (-8n*8)= b ???

OpenStudy (sachintha):

\(\Large a\times b= 0 \quad\Rightarrow \quad a=0 ~{\sf or}~ b=0\) \(\Large -8n\times (8n-8)= 0 \quad\Rightarrow \quad?\)

OpenStudy (kmeis002):

Note we said -8n and (8n-8) are multiplying, not the ones you have shown. It seems you've used the distributive property, but this undoes the multiplication of the factors.

OpenStudy (blurryface):

clearly, I'm confused.. If i'm asking you guys for help, how should I know what to do?

OpenStudy (kmeis002):

Yes, but we are concerned about the disconnect. The answer should be a = -8n and b = (8n-8) because these are the factors being multiplied. By distributing, the problem no longer has two factors being multiplied equal to zero.

OpenStudy (blurryface):

-8n=0 and 8n-8 =0

OpenStudy (kmeis002):

Correct, then just solve for n

OpenStudy (blurryface):

I already know the answer is (1,-2)...

OpenStudy (kmeis002):

Where was that gotten from? It should be 0, 1.

OpenStudy (blurryface):

I'm here to learn, not just to take the final answer and leave.. btw yea I meant (0,1)

OpenStudy (blurryface):

So can one of you guys show me how to work this problem out?

OpenStudy (blurryface):

hello?

OpenStudy (sachintha):

\(-8n=0\) What do you get by dividing both sides by -8?

OpenStudy (usukidoll):

you have two cases -8n=0 and 8n-8=0 solve each equation for n

OpenStudy (blurryface):

n=(0/-8) n=0

OpenStudy (usukidoll):

yes n =0 for one case now solve 8n-8=0

OpenStudy (blurryface):

(8n/8) = (8/8) n=1

OpenStudy (usukidoll):

mhm so the final answer is n = 0 and n = 1

OpenStudy (usukidoll):

my real question is why did this take an hour? :/

OpenStudy (blurryface):

I'm not giving anyone a medal, this took a freaking hour.. and I solved the question myself

OpenStudy (usukidoll):

ik. I feel your pain. This should've been 2 minutes...tops

OpenStudy (blurryface):

I already knew the correct answer in the first place, I just needed somebody to help show me the steps.